EX 11-13 Payroll entries Widmer Company had gross wages of $240,000 during the weck ended June 17. The amount of wages subject to social security tax was $240,000, while the amount of wages subject to federal and state unemployment taxes was $35,000. Tax rates are as follows OBJ.3 Social security Medicate State unemployment Federal unemployment 60% 15% 08% The total amount withhekd from employee wages for federal taxes was $48,000. a. Journalize the entry to record the payroll for the week of June 17. b. Journalize the entry to record the payroll tax expense incurred for the week of June 17
Answer #1
a
Wages Expense 240000
        Social Security Tax Payable 14400 =240000*6%
        Medicare Tax Payable 3600 =240000*1.5%
        Employees Federal Income Tax Payable 48000
       Wages Payable 174000
b
Payroll Tax Expense 20170
        Social Security Tax Payable 14400
        Medicare Tax Payable 3600
        State Unemployment Tax Payable 1890 =35000*5.4%
       Federal Unemployment Tax Payable 280 =35000*0.8%